# Mohammad Mehdi Morovati    Florian Tambon    Mina Taraghi    Amin Nikanjam    Foutse Khomh

### v2

@@ 5 Threats to Validity
Internal validity. As users do not necessarily provide suitable tags for their questions, our search might have missed some DRL challenges. For example, post 37973108 is related to DRL, but it does not have any specific tag mentioning DRL. Nonetheless, tag usage was necessary for a consistent methodology and we believe that the number of posts gathered and analyzed (927) is sizable enough to provide a good representation of the challenges faced by DRL developers. Moreover, we used a snowballing approach to expand our basic set of DRL-related tags used to extract DRL-related SO posts, similar to previous studies ayman2019smart . In addition to extracting posts based on the DRL-related tags, we used a set of keywords to extract DRL-related SO posts without any DRL-related tags to address this issue, as other researchers followed a similar methodology peruma2022refactor .
Another source of threat to the validity of this study arises from the potential overlap between users posing questions in SO posts and the developers associated with the DRL-related GitHub repositories that are used for our survey. To address this concern, we provided a detailed description for each category and subcategory of the taxonomy within the survey, intentionally avoiding any specific information references to any SO post as an example of challenges. Despite these precautions, the prospect of overlap between these two groups of DRL developers remains a possibility. Although utilizing the duration to receive an accepted answer has been employed in several studies alshangiti2019developing ; decan2019empirical as a metric to gauge the difficulty level of SO posts, it may present a potential threat to the internal validity of this research. Reproducibility of issues in ML-enabled systems poses significant challenges and sometimes needs more time compared to traditional software systems shah2024towards . In other words, addressing certain questions in ML-enabled systems (such as those related to parallel processing) requires specific configurations or environments, so the question may be quite simple, but it requires a long time to receive an accepted answer. To address this concern, we also considered the number of posts, which has been used as an indicator to measure the difficulty level of SO posts in a few studies bangash2019developers ; hamidi2021towards . However, as discussed in Section 3, this metric yielded similar results to the methodology we employed (i.e., the time to receive an accepted answer).
@@ 6 Conclusion and Future Works
In this study, we conducted a large-scale empirical study of 927 DRL-related posts extracted from SO. We examined all posts manually to identify the challenges that developers face when developing DRL applications. We found that Python is by far the most popular programming language and TensorFlow, OpenAI Gym, and Keras are the most frequently used libraries/frameworks for developing DRL applications. We categorized DRL development challenges into five groups including DRL issues, parallel processing & multi-threading, DRL libraries/frameworks, DL issues, and general programming. An analysis of the response received by posts reporting about the challenges shows that DRL comprehension, DRL libraries/frameworks API usage, and designing a problem using DRL algorithms are the most challenging parts of DRL applications’ development. Furthermore, parallel processing/multi-threading and DRL libraries/frameworks challenges required a longer time before receiving an accepted answer. We proposed a taxonomy of challenges and validated it using a survey of 65 DRL developers. The developers confirmed the frequency, severity, and required effort to address identified challenges. We hope that the results reported in this paper will stimulate the development of DRL quality assurance tools and guide the research community toward solving the identified challenges.

### v3

@@ 5 Threats to Validity
Internal validity. As users do not necessarily provide suitable tags for their questions, our search might have missed some DRL challenges. For example, post 37973108 is related to DRL, but it does not have any specific tag mentioning DRL. Nonetheless, tag usage was necessary for a consistent methodology and we believe that the number of posts gathered and analyzed (927) is sizable enough to provide a good representation of the challenges faced by DRL developers. Moreover, we used a snowballing approach to expand our basic set of DRL-related tags used to extract DRL-related SO posts, similar to previous studies ayman2019smart . In addition to extracting posts based on the DRL-related tags, we used a set of keywords to extract DRL-related SO posts without any DRL-related tags to address this issue, as other researchers followed a similar methodology peruma2022refactor . Another source of threat to the validity of this study arises from the potential overlap between users posing questions in SO posts and the developers associated with the DRL-related GitHub repositories that are used for our survey. To address this concern, we provided a detailed description for each category and subcategory of the taxonomy within the survey, intentionally avoiding any specific information references to any SO post as an example of challenges. Despite these precautions, the prospect of overlap between these two groups of DRL developers remains a possibility. Although utilizing the duration to receive an accepted answer has been employed in several studies alshangiti2019developing ; decan2019empirical as a metric to gauge the difficulty level of SO posts, it may present a potential threat to the internal validity of this research. Reproducibility of issues in ML-enabled systems poses significant challenges and sometimes needs more time compared to traditional software systems shah2024towards . In other words, addressing certain questions in ML-enabled systems (such as those related to parallel processing) requires specific configurations or environments, so the question may be quite simple, but it requires a long time to receive an accepted answer. To address this concern, we also considered the number of posts, which has been used as an indicator to measure the difficulty level of SO posts in a few studies bangash2019developers ; hamidi2021towards . However, as discussed in Section 3, this metric yielded similar results to the methodology we employed (i.e., the time to receive an accepted answer).
@@ 6 Conclusion and Future Works
In this study, we conducted a large-scale empirical study of 927 DRL-related posts extracted from SO. We examined all posts manually to identify the challenges that developers face when developing DRL applications. We found that Python is by far the most popular programming language and TensorFlow, Keras, PyTorch, and OpenAI Gym are the most frequently used libraries/frameworks for developing DRL applications. We categorized DRL development challenges into five groups including DRL issues, parallel processing & multi-threading, DRL libraries/frameworks, DL issues, and general programming. An analysis of the received response by the investigated SO posts shows that DRL comprehension, DRL libraries/frameworks API usage, and designing a problem using DRL algorithms are the most challenging parts of DRL applications’ development. Furthermore, parallel processing/multi-threading and DRL libraries/frameworks challenges required a longer time to receive an accepted answer. We proposed a taxonomy of challenges and validated it using a survey of 65 DRL developers. The developers confirmed the frequency, severity, and required effort to address identified challenges. We hope that the reported results in this paper will stimulate the development of DRL quality assurance tools and guide the research community toward solving the identified challenges.